OpenStudy (anonymous):

9. Solve. 6 – 3x = 3(8 – x) (Points : 4) –4 4 no solutions all real numbers can i please have help with this one, i cant get it

OpenStudy (anonymous):

what unit is this?

OpenStudy (anonymous):

9

OpenStudy (anonymous):

move the x in one side

OpenStudy (anonymous):

i know the -4 or 4 are matching up i just dont know if it is all real numbers or no solution

OpenStudy (anonymous):

6-3x=3(8-x) 6-3x=24-3x adding +3x to both side 6-3x+3x=24-3x+3x 6=24 no solution

OpenStudy (anonymous):

Thank you!

OpenStudy (anonymous):

YW